# Discrete-to-continuum limits of semilinear stochastic evolution equations in Banach spaces

### v1

@@ Abstract.
We study the convergence of semilinear parabolic stochastic evolution equations, posed on a sequence of Banach spaces approximating a limiting space and driven by additive white noise projected onto the former spaces. Under appropriate uniformity and convergence conditions on the linear operators, nonlinear drifts and initial data, we establish convergence of the associated mild solution processes when lifted to a common state space. Our framework is applied to the case where the limiting problem is a stochastic partial differential equation whose linear part is a generalized Whittle–Matérn operator on a manifold $\mathcal{M}$, discretized by a sequence of graphs constructed from a point cloud. In this setting, we obtain discrete-to-continuum convergence of solutions lifted to the spaces $L^{q}(\mathcal{M})$ for $q\in[2,\infty]$.
@@ 1.1. Background and motivation
We consider the convergence of a sequence of abstract continuous-time equations, each posed in a different Banach space in order to model the approximation of an evolution SPDE by equations that are continuous in time and discrete in space. This framework covers a typical setting where the spatial domains are finite graphs and the limiting differential operator in space is approximated by the corresponding graphical variants. If the finite graphs approximate an underlying manifold, then the graphical differential operators are related to a finite-difference approximation of the differential operator. In this setting, we study the discrete-to-continuum limit of the semidiscrete SPDEs to the corresponding continuum SPDE. To show the significance of results of this form, we now discuss a few examples of semidiscrete SPDEs as well as their continuum limits.
@@
Stochastic PDEs on graphs also appear naturally as models in the physical sciences, e.g., for interacting particle systems [17] or to represent disordered media [44]. In the former case, the continuum limit represents the large particle limit in the interacting particle system.
@@ 1.2. Main results
We will now summarize the abstract setting and main discrete-to-continuum convergence results from Sections 46, which will already be applied to a graph discretization example in Section 3.
@@
  • $-A_{n}$ is a linear operator which generates a bounded analytic semigroup of bounded linear operators on $E_{n}$ or $B_{n}$;
@@
  • $(W_{n}(t))_{t\geq 0}$ is the projection onto an appropriate subspace of a cylindrical Wiener process $(W(t))_{t\geq 0}$ on a real and separable Hilbert space $H$ (more details are specified in Theorem 1.1 below);
@@
The precise assumptions (in particular, whether the operators and initial data are $E_{n}$-valued or $B_{n}$-valued) vary throughout Sections 46, see Table 1 for an overview. Depending on the setting, the mild solutions to (1.1) are either well-defined on the whole of $[0,T]$ almost surely or cease to exist at a time $t<T$ with nonzero probability; such solutions are said to be global or local, respectively.
@@
Again we refer to Table 1 for an overview of the different settings and types of solutions, with references to the precise formulations of the corresponding assumptions. The following theorem is a summary of the discrete-to-continuum approximation results for solutions to the abstract equations (1.1) in these respective settings.
@@
    If the solutions are global and $p>1$, then for all $p^{-}<p$ we have
@@
    If the solutions are local, with associated explosion times $\sigma_{n}\colon\Omega\to(0,T]$ (see (5.41) for a definition), then we have
@@ 1.3. Contributions
The abstract discrete-to-continuum approximation theorems for stochastic semilinear parabolic evolution equations driven by additive cylindrical Wiener noise—summarized in Theorem 1.1 and proved in Sections 46—complement the results from [55, 56], which establish the continuous dependence on the coefficients of semilinear equations driven by multiplicative noise in a UMD state space, i.e., convergence in the case where $E_{n}=E$ (and $B_{n}=B$) for all $n\in\overline{\mathbb{N}}$. Given the motivating applications and our aim to provide a self-contained exposition of the proofs, we make the simplifying assumptions that the UMD spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ have type 2 (which was also assumed for $E$ in [56] but not in [55]) and the driving noise is additive.
@@ 1.4. Outline
The remainder of this article is structured as follows.
In Section 2, we establish some notational conventions, collect preliminary results regarding the (deterministic) discrete-to-continuum Trotter–Kato approximation theorem, and recall the definition of stochastic integration in Banach spaces which possess the geometric properties of unconditional martingale differences (UMD) and (Rademacher) type 2. Before proceeding to state and prove the abstract discrete-to-continuum convergence results as summarized by Theorem 1.1, we demonstrate in Section 3 how they can be applied to graph discretizations of stochastic parabolic evolution equations whose linear part is a generalized Whittle–Matérn operator on a manifold. In Section 4, we consider the linear $E_{n}$-valued version of (1.1), whose solutions are also known as infinite-dimensional Ornstein–Uhlenbeck processes. These results are extended in Section 5 to allow for semilinear $E_{n}$-valued drift operators satisfying additional (local or global) Lipschitz-continuity and boundedness assumptions. In Section 6 we first treat the analogous results in the semilinear $B_{n}$-valued setting, and then establish global well-posedness and convergence for dissipative drifts with the main aim of allowing for the polynomial type nonlinearities considered in parts of Section 3. A detailed overview of the settings and assumptions considered in the various subsections of Sections 46 can be found in Table 1. Finally, in Section 7 we discuss some potential directions for further research.
This work is complemented by three appendices: Appendix A consists of proofs of some intermediate results from Section 3 which are postponed for the sake of readability. Appendix B contains results regarding fractional parabolic integration operators which are used in Sections 4, 5.1 and 6.1. Finally, Appendix C lists the definition of (uniformly) sectorial linear operators, and some if its consequences which are used in Section 4. See Figure 1 for a schematic overview of the relations between the main Sections 36 of the manuscript and Appendices AC.
@@ 2.1. Notation
$\mathscr{L}(E)$ abbreviation for $\mathscr{L}(E;E)$
@@ 2.2. Discrete-to-continuum Trotter–Kato approximation theorem
2.3. Stochastic integration in UMD type 2 Banach spaces
@@
In order to extend the definition of the stochastic integral beyond elementary integrands, one needs to impose further geometric assumptions on the Banach space $E$. In this article we work in one of the standard settings, namely that of spaces with unconditional martingale differences and Rademacher type $2$ (abbreviated to UMD type 2). Definitions of these notions can be found in [48, Section 4.2] and [49, Section 7.1], respectively, but we will only use them to ensure existence of stochastic integrals. In this case, one can establish the following form of the Itô inequality
$\bigg{\lVert}\int_{0}^{\infty}\Phi(t)\mathop{}\!\mathrm{d}W(t)\bigg{\rVert}_{L^{2}(\Omega;E)}\lesssim_{E}\|\Phi\|_{L^{2}(0,\infty;\gamma(H;E))},$ (2.7)
@@ 3. Graph-based semilinear stochastic evolution equations with Whittle–Matérn linear operators
Before developing the general discrete-to-continuum convergence results summarized by Theorem 1.1 in the upcoming sections, in this section we demonstrate how they can be applied to the particular case of equations whose linear parts are graph discretizations of a generalized Whittle–Matérn operator on a manifold. In the spatial and linear case, such convergence results have been proven in [71, 72]. We also mention the work [65], in which the statistical properties of the spatiotemporal linear equation was investigated for fixed $n\in\mathbb{N}$.
@@ 3.1. Geometric graphs and generalized Whittle–Matérn operators
Suppose that $\mathcal{M}$ is an $m$-dimensional smooth, connected, compact manifold without boundary, embedded in $\mathbb{R}^{d}$ for some $m,d\in\mathbb{N}$. Let $\mu$ and $d_{\mathcal{M}}$ denote the uniform probability measure and geodesic metric on $\mathcal{M}$, respectively.
@@
Later on, we will need $q\in[2,\infty)$, so that $E_{n}$ is a UMD type 2 space for use in stochastic integration, but the statements here hold for all $q\in[1,\infty]$.
@@
Since $\operatorname{Id}_{n}+\mathcal{L}_{n}^{\tau,\kappa}$ is self-adjoint, positive definite and has a compact inverse for all $n\in\overline{\mathbb{N}}$ (cf. [75, Chapter XII] in the case $n=\infty$), there exists an orthonormal basis $(\psi^{(j)}_{n})_{j=1}^{n}$ of $L^{2}(\mathcal{M}_{n})$ and a non-decreasing sequence $(\lambda^{(j)}_{n})_{j=1}^{n}\subseteq[0,\infty)$, accumulating only at infinity for $n=\infty$, such that $\mathcal{L}_{n}^{\tau,\kappa}\psi^{(j)}_{n}=\lambda^{(j)}_{n}\psi^{(j)}_{n}$ for all $j\in\{1,\dots,n\}$. We summarize this state of affairs by saying that $(\psi^{(j)}_{n},\lambda^{(j)}_{n})_{j=1}^{n}$ is an orthonormal eigenbasis of $L^{2}(\mathcal{M}_{n})$ associated to $\mathcal{L}_{n}^{\tau,\kappa}$. The asymptotic behavior of the eigenvalues $(\lambda_{\infty}^{(j)})_{j\in\mathbb{N}}$ are described by Weyl’s law, cf. [75, Theorem XII.2.1]:
@@
$\Sigma_{\eta}:=\{\lambda\in\mathbb{C}\setminus\{0\}:\arg\lambda\in(-\eta,\eta)\}$ (3.14)
@@ 3.2. Convergence of graph-discretized semilinear SPDEs
    The nonlinearities $(f_{n})_{n\in\overline{\mathbb{N}}}$ are globally Lipschitz continuous and of linear growth, both uniformly in $n$. I.e., there exist $\widetilde{L}_{f},\widetilde{C}_{f}\in[0,\infty)$ such that, for all $n\in\overline{\mathbb{N}}$ and $x,y\in\mathbb{R}$,
@@
One example of a nonlinearity of the form asserted in Assumption 3.7(ii) is the cubic polynomial $f_{n}(\omega,t,x)\coloneqq-x^{3}+x$, which turns (3.16) into the stochastic Allen–Cahn equation in case $\tau=0$, $\kappa=1$ and $s=1$. Moreover, this is an example of the important situation where (3.21) is trivially satisfied by taking the same function $f_{n}\coloneqq f$ for all $n\in\overline{\mathbb{N}}$.
@@
The interplay of the various choices of spatial domains $\mathcal{M}_{n}$, linear operators $A_{n}$ and nonlinearity functions $f_{n}$ determines the class of SPDEs to which (3.16) belongs. Rigorous definitions of the corresponding mild solution concepts, as well as well-posedness and discrete-to-continuum convergence results can be found in Sections 46, respectively. Applying these results in their respective regimes of applicability, we derive the following discrete-to-continuum convergence results for the solutions to (3.16):
@@
    In Setting 3.3, let $\delta>0$ be such that $\frac{m}{1-\delta}<m+4+\delta$, suppose $s>m+4+\delta$ and Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$. Let Assumptions 3.4(ii), 3.6(i), 3.7(i) and 3.8 be satisfied. If $p\in[1,\infty)$ is such that $\Lambda_{n}\xi_{n}\to\xi_{\infty}$ in $L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};L^{\infty}(\mathcal{M}))$, then there exists a unique global mild solution $u_{n}$ in $L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{M}_{n})))$ to (3.16) for every $n\in\mathbb{N}$, as well as $u_{\infty}$ in $L^{p}(\Omega;C([0,T];C(\mathcal{M})))$ for $n=\infty$, and as $n\to\infty$ we have
@@
The proof, the intermediate results on which it relies, as well as the motivations behind the various assumptions listed above and their role in Theorem 3.9, will be discussed in the remaining subsections.
@@ 3.3. Intermediate results
The necessary convergence of the linear operators, given by (A3) and (A3-B), will ultimately be derived from the spectral convergence of $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\mathbb{N}}$ to $\mathcal{L}_{\infty}^{\tau,\kappa}$, i.e., the convergence of the respective eigenvalues and (lifted) eigenfunctions. In the square grid Setting 3.2, we can argue directly using closed-form expressions of all the eigenvalues and eigenfunctions involved, see Lemma 3.10 below. A subtlety arising in the random graph Setting 3.3 is that, for any $n\in\mathbb{N}$, we cannot in general control the errors $\lvert\lambda_{n}^{(j)}-\lambda_{\infty}^{(j)}\rvert$ and $\|\psi_{n}^{(j)}\circ T_{n}-\psi_{\infty}^{(j)}\|_{L^{q}(\mathcal{M})}$ for all $j\in\{1,\dots,n\}$, but only for indices $j$ up to a sufficiently small integer $k_{n}$. We present the precise statements below: Theorems 3.11 and 3.12a, which cover eigenvalue convergence and $L^{2}(\mathcal{M})$-convergence of eigenfunctions, are respectively taken from [71, Theorems 4.6 and 4.7]. Theorem 3.12b, concerning the $L^{\infty}(\mathcal{M})$-convergence of Laplacian eigenvalues, is a consequence of the main results from [12], as shown in [72, Lemma 15] and the discussion preceding it.
@@
    Suppose that $\mathcal{M}$ and its discretizations $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ are as in Setting 3.3, Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\infty)$, and Assumptions 3.4(ii) and 3.8 are satisfied. Then we have for all $\beta^{\prime}\in[\beta,\infty)$:
@@
The following property, called the uniform ultracontractivity of the semigroups $(S_{n})_{n\in\overline{\mathbb{N}}}$, will be needed in order to obtain the $L^{\infty}(\mathcal{M})$-convergence in Theorem 3.9b and c. Its proof relies on Riesz–Thorin interpolation, Assumption 3.6, and some arguments from Theorem 3.13.
@@
    In Setting 3.3, Assumption 3.5(i) or (ii) holds with corresponding $\beta$, as well as Assumptions 3.4(i), 3.6(i) and 3.8.

### v2

@@ Abstract.
We study the convergence of semilinear parabolic stochastic evolution equations, posed on a sequence of Banach spaces approximating a limiting space and driven by additive white noise projected onto the former spaces. Under appropriate uniformity and convergence conditions on the linear operators, nonlinear drifts and initial data, we establish convergence of the associated mild solution processes when lifted to a common state space. Our framework is applied to the case where the limiting problem is a stochastic partial differential equation whose linear part is a generalized Whittle–Matérn operator on a manifold $\mathcal{M}$, discretized by a sequence of graphs constructed from a (random) point cloud. In this setting we obtain various discrete-to-continuum convergence results for solutions lifted to $L^{q}(\mathcal{M})$ for $q\in[2,\infty]$, one of which recovers the $L^{\infty}$-convergence of a finite-difference discretization of certain (fractional) stochastic Allen–Cahn equations.
@@ 1.1. Background and motivation
We consider the convergence of a sequence of abstract continuous-time equations, each posed on a different Banach space in order to model the approximation of an evolution SPDE by equations that are continuous in time and discrete in space. This framework covers a typical setting where the spatial domains are finite graphs and the limiting differential operator in space is approximated by the corresponding graphical variants. If the finite graphs approximate an underlying manifold, then the graphical differential operators are related to a finite-difference approximation of the differential operator. As an example, we study a class of semilinear SPDEs whose linear part is given by a fractional-order elliptic differential operator on a manifold (known as a Whittle–Matérn operator), discretized by a sequence of graphs constructed from a (possibly randomly sampled) cloud of points. Linear equations of this type have previously been studied in the context of statistics and machine learning [65, 71, 72]. By verifying in detail that the hypotheses of our abstract framework are satisfied in this situation, we establish the discrete-to-continuum convergence of this scheme. Although the main advantage of our general results is their applicability to highly unstructured discretizations, we additionally show that they recover the $L^{\infty}$-convergence of finite-difference discretizations of the fractional stochastic Allen–Cahn equation on the one-dimensional flat torus. To further illustrate the significance of results of this form, we now discuss a few examples of semidiscrete SPDEs as well as their continuum limits.
@@
Stochastic PDEs on graphs also appear naturally as models in the physical sciences, e.g., for interacting particle systems [17] or the representation of disordered media [44]. In the former case, the continuum limit represents the large particle limit in the interacting particle system.
@@ 1.2. Main results
We will now summarize the abstract setting and main discrete-to-continuum convergence results from Sections 46, which will already be applied in Section 3 to the Whittle–Matérn and stochastic fractional Allen–Cahn equations described in Section 1.1.
@@
  • $-A_{n}$ is the generator of a bounded analytic semigroup of bounded linear operators on $E_{n}$ or $B_{n}$;
@@
  • $(W_{n}(t))_{t\geq 0}$ is the projection, onto an appropriate subspace, of an $H$-valued cylindrical Wiener process $(W(t))_{t\geq 0}$ for a real and separable Hilbert space $H$ (more details are specified in Theorem 1.1 below);
@@
The precise assumptions (in particular, whether the operators and initial data are $E_{n}$-valued or $B_{n}$-valued) vary throughout Sections 36; an overview is provided in Table 1. Depending on the setting, the mild solutions to (1.1) are either well defined on the whole of $[0,T]$ almost surely, or cease to exist at a time $t<T$ with nonzero probability; such solutions are said to be global or local, respectively.
@@
Again we refer to Table 1 for an overview of the different settings and types of solutions, with references to the precise formulations of the corresponding assumptions; the setting in the first row (i.e., Section 3.2) covers the (fractional) stochastic Allen–Cahn equations announced in Section 1.1, see Example 3.8 below. The following theorem is a summary of the discrete-to-continuum approximation results for solutions to the abstract equations (1.1) in these respective settings.
@@
    If the solutions are global and $p>1$, then for all $p^{-}\in[1,p)$ we have
@@
    If the solutions are local, with associated explosion times $\sigma_{n}\colon\Omega\to(0,T]$ (precisely defined in (5.41) below), then we have
@@ 1.3. Contributions
The abstract discrete-to-continuum approximation theorems for stochastic semilinear parabolic evolution equations driven by additive cylindrical Wiener noise—summarized in Theorem 1.1 and proved in Sections 46—complement the results from [55, 56], which establish the continuous dependence on the coefficients of semilinear equations driven by multiplicative noise in a state space with unconditional martingale differences (UMD), i.e., convergence in the case where $E_{n}=E$ (and $B_{n}=B$) for all $n\in\overline{\mathbb{N}}$. Given the motivating applications and our aim to provide a self-contained exposition of the proofs, we make the simplifying assumptions that the UMD spaces $(E_{n})_{n\in\overline{\mathbb{N}}}$ have Rademacher type 2 (which was also assumed for $E$ in [56] but not in [55]) and the noise is additive.
@@ 1.4. Outline
The remainder of this article is structured as follows. In Section 2, we establish some notational conventions and collect preliminaries regarding the (deterministic) discrete-to-continuum Trotter–Kato approximation theorem and stochastic integration in UMD-type-2 Banach spaces. We demonstrate in Section 3 how the results summarized by Theorem 1.1 can be applied to graph discretizations of stochastic parabolic evolution equations whose linear part is a generalized Whittle–Matérn operator on a manifold. In Section 4, we consider the linear $E_{n}$-valued version of (1.1), whose solutions are also known as infinite-dimensional Ornstein–Uhlenbeck processes. These results are extended in Section 5 to allow for semilinear $E_{n}$-valued drift operators under (local or global) Lipschitz continuity and boundedness assumptions. In Section 6 we first treat the analogous results in the semilinear $B_{n}$-valued setting, and then establish global well-posedness and convergence for dissipative drifts. Finally, in Section 7 we discuss some potential directions for further research. This work is complemented by three appendices: Appendix A consists of postponed proofs of some intermediate results from Section 3. Appendices B and C are concerned with fractional parabolic integration and (uniformly) sectorial linear operators, respectively. See Figure 1 for a schematic overview of the relations between Sections 36 and the appendices.
@@ 2.1. Notation
$\mathscr{L}(E)$ abbreviation of $\mathscr{L}(E;E)$
@@ 2.2. Discrete-to-continuum Trotter–Kato approximation theorem
2.3. Stochastic integration in UMD-type-2 Banach spaces
@@
In order to extend the definition of the stochastic integral beyond elementary integrands, one needs to impose further geometric assumptions on the Banach space $E$. In this article we work in one of the standard settings, namely that of spaces with unconditional martingale differences and Rademacher type $2$ (abbreviated to UMD-type-2). Definitions of these notions can be found in [48, Section 4.2] and [49, Section 7.1], respectively, but we will only use them to ensure existence of stochastic integrals. In this case, one can establish the Itô inequality
$\bigg{\lVert}\int_{0}^{\infty}\Phi(t)\mathop{}\!\mathrm{d}W(t)\bigg{\rVert}_{L^{2}(\Omega;E)}\lesssim_{E}\|\Phi\|_{L^{2}(0,\infty;\gamma(H;E))}$ (2.7)
@@ 3. Graph-based semilinear stochastic evolution equations with Whittle–Matérn linear operators
Before developing the general discrete-to-continuum convergence results summarized by Theorem 1.1 in the upcoming sections, in this section we demonstrate how they can be applied to the particular case of equations whose linear parts are graph discretizations of a generalized Whittle–Matérn operator on a manifold. In the spatial and linear case, such convergence results have been proven in [71, 72]. We also mention the work [65], in which the statistical properties of the spatiotemporal linear equation were investigated for fixed $n\in\mathbb{N}$.
@@ 3.1. Geometric graphs and generalized Whittle–Matérn operators
Given $m,d\in\mathbb{N}$, suppose that $\mathcal{M}$ is an $m$-dimensional smooth, connected, compact Riemannian manifold without boundary, embedded smoothly and isometrically into $\mathbb{R}^{d}$. Let $\mu$ and $d_{\mathcal{M}}$ denote the normalized volume measure and geodesic metric on $\mathcal{M}$, respectively.
@@
Later on, we will need $q\in[2,\infty)$, so that $E_{n}$ is a UMD-type-2 space for use in stochastic integration, but the statements here hold for all $q\in[1,\infty]$.
@@
Since $\operatorname{Id}_{n}+\mathcal{L}_{n}^{\tau,\kappa}$ is self-adjoint, positive definite and has a compact inverse for all $n\in\overline{\mathbb{N}}$ (cf. [75, Chapter XII] in the case $n=\infty$), there exists an orthonormal basis $(\psi^{(j)}_{n})_{j=1}^{n}$ of $L^{2}(\mathcal{M}_{n})$ and a non-decreasing sequence $(\lambda^{(j)}_{n})_{j=1}^{n}\subseteq[0,\infty)$, accumulating only at infinity for $n=\infty$, such that $\mathcal{L}_{n}^{\tau,\kappa}\psi^{(j)}_{n}=\lambda^{(j)}_{n}\psi^{(j)}_{n}$ for all $j\in\{1,\dots,n\}$. We summarize this state of affairs by saying that $(\psi^{(j)}_{n},\lambda^{(j)}_{n})_{j=1}^{n}$ is an orthonormal eigenbasis of $L^{2}(\mathcal{M}_{n})$ associated to $\mathcal{L}_{n}^{\tau,\kappa}$. The asymptotic behavior of the eigenvalues $(\lambda_{\infty}^{(j)})_{j\in\mathbb{N}}$ is described by Weyl’s law, cf. [75, Theorem XII.2.1]:
@@
$\Sigma_{\eta}\coloneqq\{\lambda\in\mathbb{C}\setminus\{0\}:\arg\lambda\in(-\eta,\eta)\}$ (3.14)
@@ 3.2. Convergence of graph-discretized semilinear SPDEs
    The nonlinearities $(f_{n})_{n\in\overline{\mathbb{N}}}$ are globally Lipschitz continuous and grow linearly, both uniformly in $n$. I.e., there exist $\widetilde{L}_{f},\widetilde{C}_{f}\in[0,\infty)$ such that, for all $n\in\overline{\mathbb{N}}$ and $x,y\in\mathbb{R}$,
@@
Example 3.8.
The cubic polynomial $f_{n}(\omega,t,x)\coloneqq-x^{3}+x$, which turns (3.16) into the (fractional) stochastic Allen–Cahn equation in case $\tau\equiv 0$, $\kappa\equiv 1$ and $s\leq 1$, is of the form asserted in Assumption 3.7(ii). Note that this is also an example of the important situation where (3.21) is trivially satisfied by taking the same function $f_{n}\coloneqq f$ for all $n\in\overline{\mathbb{N}}$.
@@
The interplay of the various choices of spatial domains $\mathcal{M}_{n}$, linear operators $A_{n}$ and nonlinearity functions $f_{n}$ determines the class of SPDEs to which (3.16) belongs. Rigorous definitions of the corresponding mild solution concepts, as well as well-posedness and discrete-to-continuum convergence results can be found in Sections 46, respectively. Applying these results in their respective regimes of applicability yields the following discrete-to-continuum convergence theorem for the solutions to (3.16); note that the setting of part c covers the stochastic (fractional) Allen–Cahn equation on the one-dimensional torus, see Example 3.8.
@@
    In Setting 3.3, let $\delta>0$ be such that $\frac{m}{1-\delta}<m+4+\delta$, suppose $s>m+4+\delta$ and Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\frac{1}{2})$. Let Assumptions 3.4(ii), 3.6(i), 3.7(i) and 3.9 be satisfied. If $p\in[1,\infty)$ is such that $\Lambda_{n}\xi_{n}\to\xi_{\infty}$ in $L^{p}(\Omega,\mathcal{F}_{0},\mathbb{P};L^{\infty}(\mathcal{M}))$, then there exists a unique global mild solution $u_{n}$ in $L^{p}(\Omega;C([0,T];L^{\infty}(\mathcal{M}_{n})))$ to (3.16) for every $n\in\mathbb{N}$, as well as $u_{\infty}$ in $L^{p}(\Omega;C([0,T];C(\mathcal{M})))$ for $n=\infty$, and as $n\to\infty$ we have
@@
The proof is presented in Section 3.4. In the next section, we list the intermediate results on which it relies. The motivations behind the various assumptions listed above, as well as their role in Theorem 3.10, are discussed in Section 3.5.
@@ 3.3. Intermediate results
The necessary convergence of the linear operators, given by (A3) and (A3-B), will ultimately be derived from the spectral convergence of $(\mathcal{L}_{n}^{\tau,\kappa})_{n\in\mathbb{N}}$ to $\mathcal{L}_{\infty}^{\tau,\kappa}$, i.e., the convergence of the respective eigenvalues and (lifted) eigenfunctions. In the square grid Setting 3.2, we can argue directly using closed-form expressions of all the eigenvalues and eigenfunctions involved, see Lemma 3.11 below. A subtlety arising in the random graph Setting 3.3 is that, for any $n\in\mathbb{N}$, we cannot in general control the errors $\lvert\lambda_{n}^{(j)}-\lambda_{\infty}^{(j)}\rvert$ and $\|\psi_{n}^{(j)}\circ T_{n}-\psi_{\infty}^{(j)}\|_{L^{q}(\mathcal{M})}$ for all $j\in\{1,\dots,n\}$, but only for indices $j$ up to a sufficiently small integer $k_{n}$. We present the precise statements below: Theorems 3.12 and 3.13a, which cover eigenvalue convergence and $L^{2}(\mathcal{M})$-convergence of eigenfunctions, are respectively taken from [71, Theorems 4.6 and 4.7]. Theorem 3.13b, concerning the $L^{\infty}(\mathcal{M})$-convergence of Laplacian eigenvalues, is a consequence of the main results from [12], as shown in [72, Lemma 15] and the discussion preceding it.
@@
    Suppose that $\mathcal{M}$ and its discretizations $(\mathcal{M}_{n})_{n\in\mathbb{N}}$ are as in Setting 3.3, Assumption 3.5(ii) holds with $\beta\in(\frac{m+4+\delta}{2s},\infty)$, and Assumptions 3.4(ii) and 3.9 are satisfied. Then we have for all $\beta^{\prime}\in[\beta,\infty)$:
@@
The following property, which we call the uniform ultracontractivity of the semigroups $(S_{n})_{n\in\overline{\mathbb{N}}}$, will be needed in order to obtain the $L^{\infty}(\mathcal{M})$-convergence in Theorem 3.10b and c. Its proof relies on Riesz–Thorin interpolation, Assumption 3.6, and some arguments from Theorem 3.14.
@@
    In Setting 3.3, Assumption 3.5(i) or (ii) holds with corresponding $\beta$, as well as Assumptions 3.4(i), 3.6(i) and 3.9.